OpenStudy (anonymous):

Which equation is in slope-intercept form and represents a line with slope 3 through the point (9, –4)?

OpenStudy (anonymous):

what are your suggestions ?

OpenStudy (anonymous):

A. -3x+y=21 B. y=3x-31 c. y=3x+21 D. y=-3x+y=-31

OpenStudy (anonymous):

@horrible.com

OpenStudy (alexandervonhumboldt2):

B

OpenStudy (anonymous):

thank you